(sqrt(8) - 3)(sqrt(8) + 4)
Answer by algebrahouse.com(1659)

(√8 - 3)(√8 + 4)
= √8 (√8) + √8(4) - 3(√8) - 3(4) {used foil method}
= √64 + 4√8 - 3√8 - 12 {multiplied through}
= 8 + √8 - 12 {evaluated square root of 64 and combined like terms}
= -4 + √4 √2 {combined like terms and simplified the square root of 8}
= -4 + 2√2 {evaluated the square root of 4}

For more help from me, visit: www.algebrahouse.com